# LingoQA: Video Question Answering for Autonomous Driving

### v1

@@ 5.2 Training Datasets Mixture
Table 4 shows the contributions of our two datasets, Action Dataset and the Scenery Dataset. Both datasets proved influential in improving model performance. We show that fine-tuning on the LingoQA dataset that we open source leads to a considerable improvement compared to general pre-training only.
@@ 5.3 Impact of Frame Count
We want to investigate the variation in VQA performance with decreasing and increasing the number of video frames fed into the model. The base model contains 5 frames over a 4-second context. The performance declined when shifting from multi-frame video to a single image representation, which can be explained by the model not getting enough information to answer questions where temporal information is crucial. The performance when increasing the number of frames to 3, 5 and 7 remains relatively consistent. We hypothesise this is due to the lack of effective video fusion and leave it to future work to investigate video-language encoders further,
@@ 5.4 Impact of Video Fusion Strategy
Given how crucial temporal context is for scenario understanding, this study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion. The early-fusion method employs average pooling to condense features from the vision encoder prior to their incorporation into the Q-Former, producing a unified visual feature vector for language space projection. In contrast, the mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism. The late-fusion method, where individual frame embeddings from Q-Former output are fed into the LLM, allowing it to resolve temporal relationships. Our findings demonstrate that both mid-fusion and late-fusion are effective methods for incorporating video content into the model. Mid-fusion allows for a greater number of context tokens through the use of a predetermined number of video embeddings. Conversely, late-fusion shows a slightly enhanced performance by providing comprehensive frame information to the LLM.
@@ 5.5 Impact of Large Language Model
We investigate the impact that different Large Language Models have on the overall performance of our vision-language model. As shown in Table 4, the best score is achieved by the Vicuna-1.5-7B [13] that our base model uses. In the same model family Llama-2-7B [48] achieves comparable, but slightly lower performance. Mistral-7B [28], despite its promise of superior performance over Llama-2, proved less effective in our fine-tuning task. Notably, OPT-7B [60] demonstrates significantly lower performance, despite having a similar number of parameters. This discrepancy underscores the crucial role of the pretraining phase in the base language model’s effectiveness.
@@ 6 Discussion
Limitations of Lingo-Judge.
This work still has a few limitations that we discuss and provide guidelines as to how their effects may be mitigated. The primary limitation of our proposed Lingo-Judge is that we have not studied its ability to generalise to other datasets and to other domains. The classifier has been trained for autonomous driving evaluation and has been shown effective for this purpose. Hence, the classifier can be used for evaluation when paired with the benchmark dataset that we release with the paper but would need to be re-trained if employed on a new dataset. Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset. The same response style is adopted in the LingoQA training sets and the models. We leave generalization to different answering styles to future work. Notably, Lingo-Judge does rate human labels as highly truthful, providing confidence in its ability to rate typical human answers. Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer. We leave it to future work to extend the classifier to predict human preference scores directly instead of factual correctness.
@@
One of the primary constraints is that our model operates on relatively short video segments and few frames, limiting the contextual understanding of scenarios. We also do not test for driving decisions and attention mechanisms, focusing on question-answering abilities only. We did not test the scaling in our models and focused on the most practical 7B parameter LLMs only. Our dataset and baseline are limited to information from a single front-facing car camera, excluding additional sensory inputs like LiDAR that could enrich the model’s understanding of the driving environment. Expanding the model to address the short video context, as well as adding action prediction and evaluation to the dataset and the benchmark would result in a more robust and versatile system for autonomous driving.
@@ 7 Conclusion
In this paper, we introduced a novel benchmark for Video Question Answering for autonomous driving. The benchmark consists of a evaluation dataset, learned classifier-based metric Lingo-Judge that is highly correlated with human evaluation, a comprehensive high-quality training dataset for autonomous driving. The fast feedback from employing Lingo-Judge facilitates effective exploration in the video QA field. Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of explainability in end-to-end autonomous driving systems. The LingoQA benchmark is openly released to spur further community research, providing a reliable and highly correlated evaluation method to human ratings.
@@ Appendix A LingoQA Dataset Examples
Further examples on the capabilties existent in the training and the evaluation datasets are shown in Figure 4. The scenery dataset contains highly descriptive elements, such as object colours, junction type, construction zones, traffic lights, and the road layout. The action dataset is complementary and focused on driving competencies, such as the impact of traffic lights on driving and interactions with other road agents. The evaluation dataset contains a broad range of questions aimed to test competencies relevant for autonomous driving.
@@ Appendix B Lingo-Judge Examples
Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
How many pedestrians are crossing the road? Zero pedestrians There are no pedestrians crossing the road. 23.18 5 5 0.96 True
What is the road speed limit? 20 mph - it is written on the road The road speed limit is 20 mph. 97.91 4 3 0.95 True
How many cars are driving in your direction? None There are no cars driving in my direction. 0.11 4 3 0.96 True
Which vehicle should you follow if any? The motorcyclist. If any, I should follow the motorcycle ahead. 0.42 4 5 0.95 True
What is the current action and its justification? Answer in the form “action, justification” Slow down, there is a stationary van infront of us I am decelerating because of the stationary truck ahead. 34.79 5 5 0.96 True
What is the current action and its justification? Answer in the form “action, justification” Stop, Red light I am stopping because the traffic lights to go straight are red. 18.62 5 5 0.95 True
How many cyclists can you see? I can see 3 cyclists I can see two cyclists. 150.29 1 2 0.05 False
What color are the traffic lights showing? The traffic lights are showing green The traffic lights are showing red. 329.36 0 1 0.05 False
What action are you taking with respect to the cyclist? Overtaking them on the right and keeping the speed I am overtaking the cyclist on the left. 349.52 2 2 0.10 False
In which direction is the bus driving? The bus is driving in the opposite direction The bus is driving in the oncoming direction. 404.65 4 5 0.31 False
Are there any parked car on the side of the road? Yes, there are two cars parked on the right of the road No, there are no parked cars on either side of the road. 142.40 0 0 0.05 False
Is acceleration necessary in this situation? If so, provide the reason. No. We should decelerate in this situation because there is a vehicle stopping ahead of us. No, acceleration is not necessary in this situation as I am already driving at the speed limit. 177.96 3 3 0.31 False
@@
Question Label Prediction CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
Is any vehicle stopped at the zebra crossing? There is no zebra crossing. No, there are no vehicles stopped at the zebra crossing. 106.19 4 2 0.93 True
What elements are capturing your attention and what is the reason behind it? I am making sure there is no vehicle on the left lane I am making my lane change to. I am also looking at the traffic light to make sure it stays green while I am crossing the intersection. I am paying attention to the bus lane on the left because I need to maintain a safe distance from it. 0.91 3 1 0.57 True
What is the current action and its justification? Answer in the form ”action, justification” We are taking a left turn, because we are at a t-junction. Then we keep a safe distance from the cyclists in front. I am decelerating to keep a safe distance from the cyclist ahead of me. 46.45 2 2 0.32 False
@@ Appendix D Lingo-Judge Correlation Study
In Figure 8, the metric scores are plotted against the human evaluators’ grades (from 0 to 1). In red is the least-squares regression of the linear relationship between the metric and the human-assigned grades. Figure 9 shows the value of both correlation coefficients for each of the 5 metrics, as well as their confidence interval bounds. We note than not only does Lingo-Judge provided higher correlation, it also provides tighter confidence intervals than the other metrics.
@@
Appendix E Training Parameters
@@
Appendix F LingoQA Baseline Examples

### v2

@@ 5.2 Training Datasets Mixture
Table 4.2 shows the contributions of our two datasets, Action Dataset and the Scenery Dataset. Both datasets proved influential in improving model performance. We show that fine-tuning on the LingoQA dataset that we open source leads to a considerable improvement compared to general pre-training only.
@@ 5.3 Impact of Frame Count
We want to investigate the variation in VQA performance with decreasing and increasing the number of video frames fed into the model. The base model contains 5 frames over a 4-second context. The performance declined when shifting from multi-frame video to a single image representation, which can be explained by the model not getting enough information to answer questions where temporal information is crucial. The performance when increasing the number of frames to 3, 5, and 7 remains relatively consistent. We hypothesize that this is due to the lack of effective video fusion, hence we highlight the opportunity for improved video encoding methods.
@@ 5.4 Impact of Video Fusion Strategy
Given how crucial the temporal context is for scenario understanding, this study explores three methods for integrating video frames into the LLM: early-fusion, mid-fusion, and late-fusion. The early-fusion method employs average pooling to condense features from the vision encoder prior to their incorporation into the Q-Former, producing a unified visual feature vector for language space projection. In contrast, the mid-fusion approach, merges video features into fixed-size tokens within the Q-Former with the cross-attention mechanism. The late-fusion method, where individual frame embeddings from Q-Former output are fed into the LLM, allows it to resolve temporal relationships. Our findings demonstrate that both mid-fusion and late-fusion are effective methods for incorporating video content into the model. Mid-fusion allows for a greater number of context tokens through the use of a predetermined number of video embeddings. Conversely, late-fusion shows a slightly enhanced performance by providing comprehensive frame information to the LLM.
@@ 5.5 Impact of Large Language Model
We investigate the impact that different Large Language Models have on the overall performance of our vision-language model. As shown in Table 4.2, the best score is achieved by Vicuna-1.5-7B [13] that our base model uses. In the same family of models, Llama-2-7B [48] achieves comparable, but slightly lower performance. Mistral-7B [26], despite its promise of superior performance over Llama-2, proved less effective in our fine-tuning task. Notably, OPT-7B [60] demonstrates significantly lower performance, despite having a similar number of parameters. This discrepancy underscores the crucial role of the pretraining phase in the base language model’s effectiveness.
5.6 Evaluation of SOTA Vision-Language Models
To demonstrate the relevance of the newly proposed benchmark, we evaluate a series of SOTA vision-language models and compare them to human performance. The models shown in Table 5 are evaluated zero-shot. The best performing models achieve an accuracy as rated by humans of 56.67%, which is below human performance at 93.4%. This highlights the challenging nature of our proposed benchmark.
Table 5: Evaluating vision-language models zero-shot. The performance of existing vision-language models is far from human capability, highlighting the challenging nature of the benchmark.
Human Lingo-J GPT-4 BLEU METEOR CIDEr
Human 93.4 96.6 4.68 81.04 52.92 361.77
GPT-4V 56.67 59.6 3.19 6.30 12.35 42.82
LLaVA 38.97 49.4 2.45 4.23 8.38 38.39
FUYU 17.69 45.4 2.28 1.90 13.00 12.04
@@ 6 Discussion
Limitations of Lingo-Judge.
This work still has a few limitations that we discuss and provide guidelines as to how their effects may be mitigated. The classifier has been trained for autonomous driving evaluation and has been shown effective for this purpose. Hence, the classifier can be used for evaluation when paired with the benchmark dataset that we release with the paper. Second, we optimized the classifier to evaluate responses in the style provided by human annotators in the evaluation dataset. The same response style is adopted in the LingoQA training sets and the models. Further details regarding generalisation to response styles is studied in Appendix E. Third, as the classifier is only trained to predict factual correctness, it cannot discern which answer of two equally correct answers humans prefer.
@@
One of the primary constraints is that our model operates on relatively short video segments and few frames, limiting the contextual understanding of scenarios. We also do not test for driving decisions and attention mechanisms, focusing on question-answering abilities only. We did not test the scaling in our models and focused on the most practical 7B parameter LLMs only. Our dataset and baseline are limited to information from a single front-facing car camera, excluding additional sensory inputs like LiDAR that could enrich the model’s understanding of the driving environment. Expanding the model to address the short video context, as well as adding action prediction and evaluation to the dataset and the benchmark would result in a more versatile system for autonomous driving.
@@ 7 Conclusion
In this paper, we introduced a novel benchmark for Video Question Answering for autonomous driving. The benchmark consists of a evaluation dataset, learned classifier-based metric Lingo-Judge that is highly correlated with human evaluation, a comprehensive high-quality training dataset for autonomous driving. The fast feedback from employing Lingo-Judge facilitates effective exploration in the video QA field. Additionally, the comprehensive experiments on different model combinations presented in this paper can become a foundation for further improvement of end-to-end autonomous driving systems. The LingoQA benchmark is openly released to spur further community research, providing a reliable and highly correlated evaluation method to human ratings.
@@ Appendix A LingoQA Dataset Examples
Further examples on the capabilties existent in the training and the evaluation datasets are shown in Figure 4. The scenery dataset contains highly descriptive elements, such as object colours, junction type, construction zones, traffic lights, and the road layout. The action dataset is complementary and focused on driving competencies, such as the impact of traffic lights on driving and interactions with other road agents. The evaluation dataset contains a broad range of questions aimed to test competencies relevant for autonomous driving. Further examples from the evaluation benchmark are also included in the overview video accompanying the submission. Notably, any personal identifiable information, such as faces and plate ID’s, has been annonymised in the dataset.
@@ Appendix B Lingo-Judge Examples
Question
Label
Prediction
CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
How many pedestrians are crossing the road?
Zero pedestrians
There are no pedestrians crossing the road.
23.18 5 5 0.96 True
What is the road speed limit?
20 mph - it is written on the road
The road speed limit is 20 mph.
97.91 4 3 0.95 True
How many cars are driving in your direction?
None
There are no cars driving in my direction.
0.11 4 3 0.96 True
Which vehicle should you follow if any?
The motorcyclist.
If any, I should follow the motorcycle ahead.
0.42 4 5 0.95 True
What is the current action and its justification? Answer in the form “action, justification”
Slow down, there is a stationary van infront of us
I am decelerating because of the stationary truck ahead.
34.79 5 5 0.96 True
What is the current action and its justification? Answer in the form “action, justification”
Stop, Red light
I am stopping because the traffic lights to go straight are red.
18.62 5 5 0.95 True
How many cyclists can you see?
I can see 3 cyclists
I can see two cyclists.
150.29 1 2 0.05 False
What color are the traffic lights showing?
The traffic lights are showing green
The traffic lights are showing red.
329.36 0 1 0.05 False
What action are you taking with respect to the cyclist?
Overtaking them on the right and keeping the speed
I am overtaking the cyclist on the left.
349.52 2 2 0.10 False
In which direction is the bus driving?
The bus is driving in the opposite direction
The bus is driving in the oncoming direction.
404.65 4 5 0.31 False
Are there any parked car on the side of the road?
Yes, there are two cars parked on the right of the road
No, there are no parked cars on either side of the road.
142.40 0 0 0.05 False
Is acceleration necessary in this situation? If so, provide the reason.
No. We should decelerate in this situation because there is a vehicle stopping ahead of us.
No, acceleration is not necessary in this situation as I am already driving at the speed limit.
177.96 3 3 0.31 False
@@
Question
Label
Prediction
CIDEr GPT4 GPT4-CoT L-J Prob. L-J Class.
Is any vehicle stopped at the zebra crossing?
There is no zebra crossing.
No, there are no vehicles stopped at the zebra crossing.
106.19 4 2 0.93 True
What elements are capturing your attention and what is the reason behind it?
I am making sure there is no vehicle on the left lane I am making my lane change to. I am also looking at the traffic light to make sure it stays green while I am crossing the intersection.
I am paying attention to the bus lane on the left because I need to maintain a safe distance from it.
0.91 3 1 0.57 True
What is the current action and its justification? Answer in the form ”action, justification”
We are taking a left turn, because we are at a t-junction. Then we keep a safe distance from the cyclists in front.
I am decelerating to keep a safe distance from the cyclist ahead of me.
46.45 2 2 0.32 False
@@ Appendix D Lingo-Judge Correlation Study
In Figure 8, the metric scores are plotted against the human evaluators’ grades (from 0 to 1). In red is the least-squares regression of the linear relationship between the metric and the human-assigned grades. Figure 9 shows the value of both correlation coefficients for each of the 5 metrics, as well as their confidence interval bounds. We note that not only does Lingo-Judge provide higher correlation, it also provides tighter confidence intervals than the other metrics.
Appendix E Lingo-Judge Generalisation
To investigate the generalisation abilities of the Lingo-Judge, we examine the performance of the model on a range of answer styles. In particular, we evaluate vision-language models with varying architectures namely GPT4-V, LLaVA and FUYU. Table 8 shows the performance of the Lingo-Judge as measured by the validation accuracy. This highlights that the model performs the best on short, concise answers, akin to those of humans.
Table 8: Robustness to response styles. Investigation into the impact of the response style on validation accuracy.
Lingo-J Val. Acc.
LingoQA 89.50
GPT-4V
few-shot (FS) 83.27
concise (C) 81.63
unprompted (U) 83.06
incorrect (I) 89.59
LLaVA
concise (C) 81.43
unprompted (U) 78.12
FUYU 64.89
@@ Appendix E Lingo-Judge Generalisation
Appendix F Training Parameters
@@
Appendix G LingoQA Baseline Examples